# Bullet: Boosting GPU Utilization for LLM Serving via Dynamic Spatial-Temporal Orchestration

### v1

@@ 2.4. Existing Optimizations
Apart from chunked prefill-based approaches, MuxServe (Duan et al., 2024) (Figure 3c) decouples prefill and decode phases into separate processes, leveraging MPS (NVIDIA, 2025) with manually configured, fixed SM quotas to enable spatial sharing. However, this coarse-grained partitioning relinquishes execution control to the device after launching the model, which consists of hundreds of kernels. This leads to unpredictable latency despite GPU saturation (section 2.2.2). Furthermore, the inflexible SM allocation cannot adapt to dynamic workloads, compromising serving quality. Thus, neither solution fully resolves the utilization-latency tradeoff under varying serving demands.
@@ 2.5. Opportunities and Challenges
To address the inefficiencies discussed above, a concurrent yet predictable execution mechanism is urgently demanded to maximize GPU utilization in LLM serving while adhering to latency constraints, as illustrated in Figure 5.
Concurrent Schedule and Execution. Despite the concurrent execution of prefill and decode improving performance (Duan et al., 2024), maintaining latency targets requires fine-grained request monitoring and kernel scheduling based on real-time progress and system status. A layer-level SLO-aware scheduler (section 3.3) may effectively address these challenges, but must be designed to with minimal overhead and synchronization costs (section 3.5).
Optimized Resource Provision. As discussed in section 2.2.3, neither short nor long input sequences saturate GPU resources during the prefill phase. By carefully partitioning available SM resources (section 3.4), memory-bound decode kernels may efficiently co-execute with compute-intensive prefill kernels. However, the multi-dimensional nature of input parameters and dynamic inter-kernel contention create a complex optimization space. This necessitates precise execution time modeling (section 3.2) across SM allocations and efficient exploration algorithms (section 3.3) to identify optimal configurations.
SLO-aware Fast Re-configuration. In response to runtime status dynamics, including new incoming requests or potential SLO violations, resource allocation between prefill and decode phases must be re-configured. These frequent adjustments demand near zero-overhead reconfiguration to maintain efficiency. Although MPS (NVIDIA, 2025) offers context APIs to adjust different SM quotas, we explore an even more lightweight approach (section 3.4) to maintain multiple pre-configured execution states for instantaneous switching between optimized resource partitions.
Figure 5 illustrates dynamic resource scheduling in a 4-layer LLM serving scenario. Initially, ❶ prefill and decode run concurrently with ❷ resource provision at a balanced point. When request B arrives, request A’s prefill is ❸ prioritized by allocating more resources while avoiding suspending active decodes. Once request B begins its prefill phase, the scheduler ❸ dynamically adjusts quotas again to maintain the optimal configuration that satisfies both SLO requirements. Overall, this approach enables fine-grained control of task execution for dynamic workloads.
@@ 3.1. Overview
Figure 6 illustrates Bullet’s workflow for concurrent prefill and decode execution with optimal resource provision dynamically. Bullet comprises four key components around the scheduling, resource partition and execution: performance estimator, SLO-aware task scheduler, computational resource manager and concurrent execution engine. The performance estimator (section 3.2) ❶ first builds an analytical model for the served LLM to facilitate profiling and collects profile data to improve the prediction accuracy. The enhanced model precisely estimates task latency across different configurations of co-executing prefill and decode batch sizes and allocated computing resources. During runtime, SLO-aware task scheduler (section 3.3) acts as a critical coordinator to enable GPU sharing with prefill and decode tasks via spatial-temporal scheduling. At every layer-wise scheduling cycle, the scheduler ❷ proactively retrieves system status from concurrent execution engine (section 3.5), monitoring request progress and evaluating potential SLO violations through a combination of performance estimator and historical statistics. The scheduler rapidly searches the optimal resource configuration and ❸ scheduling decision that maximizes throughput while adhering to SLO requirements. Computational resource manager is ❹ triggered for lightning resource re-configuration when necessary. Finally, the prefill and decode kernels are ❺ launched concurrently on the provisioned SMs. Bullet dynamically balancing the competing demands of TTFT and TPOT maintains high utilization.
@@ 3.2. Performance Estimator
3.2.1. Performance Modeling
Accurate performance estimation is critical for SLO-aware scheduling (section 3.3) to determine optimal resource configurations efficiently. While predicting LLM operators’ latency can be achieved using the roofline model (Agrawal et al., 2024; Yuan et al., 2024), non-trivial effects arise when considering SMs are partitioned, wave quantization effects and inter-kernel contention. These factors make purely analytical approaches insufficient. Additionally, the exponential combinations of potential kernel overlaps and input dimensions render exhaustive profiling (Zhang et al., 2025a; Strati et al., 2024) impractical. We analytically model these effects, then leverage offline profiling to refine the model with minimizes required data points while maintaining accuracy.
Consider a kernel with flops of operations and access bytes of memory, we first estimate the execution time utilizing the roofline model (Agrawal et al., 2024; Yuan et al., 2024). When the kernel is restricted to use SMs on an SMs GPU, the available compute resource and memory bandwidth non-linear scales to the ratio of , as illustrated in Figure 7. Compute intensive prefill kernels are susceptible to resource restrictions, while the memory-bound decode scales super-linear with the SMs. To model such behavior, we suppose the peak compute flops and bandwidth with decaying factors and , respectively. These factors are obtained via the subsequent profiling. Similarly, as co-located prefill and decode kernels compete for compute and bandwidth resources, we denote the decaying factor as and , respectively. Finally, by introducing the ideal SM ratio of wave quantization in Equation 1, we formulate the estimation as:
3.2.2. Offline Profiling
While the model characterizes isolated kernel performance, analyzing inter-kernel contention between co-located prefill and decode tasks presents combinatorial complexity. Therefore, we focus on Transformer layer latency rather than individual kernels, namely, . Through profiling, we realize the parameters in Equation 2 to capture layer-level behavior, enabling predictable latency analysis without explicitly tracking individual kernel interactions during concurrent prefill and decode execution. The profiling records latency varied by five parameters, prefill sequence length (), decode batch size (), average context length in decode batch (), number of SMs for prefill (), and number of SMs for decode (). The process begins with isolated kernel measurements to establish and , followed by co-located execution to fit contention parameters and across prefill and decode. Since the parameter space is highly complex, we sample data at steps, enumerating and SM counts at a step of 1024, 8, 1024 and 6 while keeping within KV cache capacity. This sampling effectively reduces the number of trails to approximately 12k while preserving coverage. Finally, the augmented model interpolates unsampled configurations using derived scaling factors and . The profiling process completes within approximately two hours, with model accuracy validated on real-world workloads in section 4.5.2.
@@ 3.3.1. Scheduling Workflow
The scheduler operates independently for prefill and decode tasks within their respective concurrent execution engines (section 3.5). At each scheduling cycle, the scheduler retrieves global system status from the shared metadata buffer (section 3.5.2) to inform the decisions. Figure 8(a) demonstrates a workflow on a 12-SM GPU executing a 4-layer LLM, with layers in the prefill phase denoted as P1-P4 and a decode step marked as D1-D7. For prefill tasks, the scheduler launches a fixed number of layers (1 in the example) and synchronizes GPU to make subsequent scheduling decisions. This enables real-time perception (section 3.3.2) of prefill progress and rapid adaptation to system fluctuations. In contrast, decode tasks are dispatched as a single compounded operation via CUDA Graph (Corporation, 2025) to minimize kernel launch overhead, with the scheduler invoked before each iteration. Each scheduling decision estimates latency through performance estimator, thereby perceiving the request progress. Based on the predicted latency, section 3.3.3 describes the primary scheduling objective, which is to meet stringent SLO requirements while maximizing throughput. If the estimated latency violates SLOs with current resource configurations, the scheduler employs sophisticated rules (Algorithm 1, 2) to rapidly determine the optimal resource partition configurations. Finally, the scheduler triggers the computational resource manager to switch resources configuration, ensuring compliance with latency guarantees.
3.3.2. Execution Progress Tracking
Scheduling decisions depend on the progress of the executing kernels, and the all requests’ latency. The system state at scheduling iteration is defined as , where represents the progress of the running prefill operation, denotes the aggregated TPOT in the decode batch, and specifies the resource allocation configuration. The prefill state , in which indicates the number of executed layers, is the total number of tokens in the prefill batch, i.e., the sum of sequence lengths of batched requests. The reflects the elapsed time since prefill starts, is the queuing time for the -th request, and is the number of pending requests in the waiting queue at step . Decode batch status tracks the batch size , output token count and decode latency for request , with TPOT calculated as . Resource configuration consists of the number of SMs allocated to prefill and decode, respectively. This comprehensive state representation enables the scheduler to jointly optimize for both immediate task progress and long-term SLO compliance while accounting for resource utilization efficiency.
Algorithm 1 outlines the scheduling process. First, the algorithm monitors prefill request progress and updates the system states (line 2-6). Also, decode tasks update per-request latency metrics to compute real-time TPOT values (line 8-10). This dual-level estimation captures both immediate system status and holistic request progression. Notably, the estimator refines its accuracy by correlating deviations between predicted latency and observed execution times across scheduling cycles, using the inter-cycle duration as dynamic feedback to adjust future predictions.
3.3.3. Scheduling Space and Objectives
The primary scheduling objective is to balance between TTFT and TPOT and prioritize prefill if decode latency is not violated, which helps optimize throughput. Based on the real-time progress tracking, the scheduler explores the optimization space to determine the optimal resource allocation and the number of concurrent decode steps that co-executes with prefill operations. When the estimated TTFT and TPOT both violate the requirements (lines 10-11), the scheduler strategically reduces the number of SMs allocated to decode tasks.This reallocation accelerates prefill execution while maintaining acceptable decode performance downgrade. However, if satisfying both constraints becomes infeasible due to a high request rate (line 12-17), the system dynamically balances resource allocation to prevent excessive latency from occurring in both stages.
For instance, when TTFT violation is detected (line 14, Figure 8(a)-❶), the scheduler reduces decode SMs to accelerate prefill. The search algorithm iteratively decreases decode SMs from the current to a minimum value , while estimating delays each step to prevent future violations. The algorithm returns when a new decode SM count is found. However, if setting SMs still violates TTFT while adhering to TPOT, the system prioritizes prefill by temporarily pausing decode (❷) until the next cycle (Figure 8(a)). This indicates significant potential to improve prefill performance by temporarily borrowing resources from decode tasks. The process repeats until find an optimal configuration, balancing prefill speed while avoiding excessive decode suspension.
@@ 3.4. Computational Resource Manager
3.4.1. Computation Unit Partitioning
Current approaches for dynamic SM partition inevitably involve intensive kernel code modification (Zhu et al., 2024) to achieve precise management. Alternatively, several non-intrusive methods (Zhang et al., 2025a; Chen et al., 2022; Duan et al., 2024) use MPS context (NVIDIA, 2025) to share GPU among different processes, which are tagged with maximum SM limitations. However, simply specifying the context offers no guarantee on which SMs would be allocated for concurrent kernels, potentially aggravating contention (Olmedo et al., 2020; Jain et al., 2019). To gain better control of SM allocation, we employ an SM mask technique (Bakita and Anderson, 2023, 2024) upon the MPS. The libsmctrl_set_stream_mask API enforces kernel execution on designated SMs by configuring available SM constraints for the task queue, namely CUDA stream (Corporation, 2025). Specifically, the API supports allocation granularity of 2 SMs and support for arbitrary SM group configurations (Figure 8(b)). Therefore, by submitting tasks to the masked streams, resource partition can be realized at a fine-grained level while maintaining the benefits of multi-process GPU sharing through MPS. Despite targeting CUDA, the design can also be extended to other hardware supporting concurrent task queues and resource partition (MxGPU, 2016).
3.4.2. Instant Layer-wise Re-configuration
To achieve agile resource re-configuration layer-wise (section 3.3), Bullet creates multiple CUDA streams with different SM partitions beforehand. During execution, the task scheduler instantly selects the pre-configured stream for kernel launches, achieving efficient spatial-temporal resource sharing through rapid stream switching. Moreover, such flexible configuration enables non-stricitly isolated configurations to allow for kernels sharing intra-SM resources.
As depicted in Figure 8, a set of streams is masked with various SM counts in advance. Initially, first prefill layer and two decode iterations onto strictly partitioned streams with 6 SMs each. Then, the scheduler estimates a potential prefill timeout and ❶ instructs the second layer to launch on a stream configured with 8 SMs. In contrast, the decode task is launched with fewer SMs. In the next scheduling cycle, decode execution is ❷ temporarily delayed to accelerate TTFT without violating TPOT constraints, delaying decode to the next cycle. At the sixth decode step, the scheduler ❸ anticipates prefill completion and ❹ allocates 8 SMs to decode, enabling controlled SM sharing with prefill. This intentional resource contention facilitates a smooth transition between co-running prefill/decode and decode-only, trading marginal throughput reduction for significantly improved hardware utilization.

### v2

@@ 2.3.3. Existing Optimizations
Despite recent optimizations of chunked prefill, the technique remains intrinsically limited. PodAttention (Prabhu et al., 2025) fuses prefill and decode attention kernels to accelerate hybrid-batch attention, yet this operator-specific optimization does not mitigate the fundamental latency penalties imposed by chunking (section 2.3.2). Nanoflow (Zhu et al., 2025) designs a fixed-size pipeline that uses CUDA streams to overlap chunked attention and linear layers with careful inter-kernel synchronizations. However, the latency of chunked attention grows by successive chunks, which eventually eliminates the chances of overlapping. Apart from chunked prefill-based approaches, MuxServe (Duan et al., 2024) (Figure 4) decouples prefill and decode phases into separate processes, leveraging MPS (NVIDIA, 2025) with manually configured, fixed SM quotas to enable spatial sharing. This coarse-grained static partitioning leads to unpredictable latency despite saturating the GPU. Therefore, neither solution fully resolves the throughput-latency tradeoff under varying serving demands.
Takeaway-2. Both chunked prefill and uncoordinated spatial sharing exhibit skewed throughput–latency tradeoffs. This urges for a fine-grained yet predictable execution mechanism for optimal LLM serving performance.
@@ 2.4. Opportunities and Challenges
To address the inefficiencies discussed above, an inter-phase orchestration system is demanded to achieve optimal throughput-latency balance while maximizing GPU utilization, as illustrated in Figure 6.
Concurrent Schedule and Execution. Co-executing prefill and decode eliminates redundant KV-cache reloads and fully utilizes GPU resources, but sustaining latency targets demands real-time request monitoring and kernel scheduling that respond to execution progress and system state. A layer-level, SLO-aware scheduler must tightly control execution advance to avoid inter-phase resource competition. While LLM online serving further requires low-overhead control plane and support for rapid reconfiguration, rendering these challenges non-trivial.
Optimized Resource Provision. By carefully partitioning available SM resources, memory-bound decode kernels may efficiently co-execute with compute-intensive prefill kernels. However, LLM’s multi-dimensional nature of input parameters and dynamic inter-kernel contention creates a complex optimization space. This necessitates precise latency modeling across varying SM provisions and efficient exploration algorithms to identify optimal configurations.
SLO-aware Rapid Re-configuration. In response to runtime workload dynamics, resource allocation between prefill and decode phases must be re-configured. These frequent adjustments demand near-zero-overhead reconfiguration to maintain efficiency. Although MPS (NVIDIA, 2025) offers context APIs (Corporation, 2025) to adjust different SM quotas, it incurs significant memory overhead when applied to LLM serving scenarios due to dynamic inputs. An even more lightweight approach is demanded to maintain for instantaneous switching between optimized resource partitions.
@@ 3.1. Overview
Figure 7 illustrates Bullet’s workflow for concurrent prefill-decode execution with dynamic and fine-grained resource provisioning, with all components operating at microsecond-level overhead. Bullet comprises four key components around the scheduling, resource partitioning and execution: performance estimator, SLO-aware task scheduler, resource manager, and concurrent execution engine. The performance estimator (section 3.2) ❶ first builds an analytical model for the served LLM, augmented with lightweight offline samples. The model provides precise latency predictions across different configurations of co-executing prefill and decode batch sizes under varying resource allocations. At runtime, the SLO-aware task scheduler (section 3.3) acts as the central coordinator to enable GPU sharing between prefill and decode tasks via spatial-temporal scheduling. During every layer-wise scheduling cycle, the scheduler ❷ proactively retrieves system status from the concurrent execution engine (section 3.5), monitoring request progress and ❸ evaluating potential SLO violations by performance estimator. Observed statistics are also used to refine the performance estimator online. The scheduler rapidly searches for ❹ an optimal resource configuration and scheduling decision that maximizes throughput while ensuring SLO compliance. Computational resource manager is ❺ triggered for lightning resource reconfiguration when necessary. Finally, the prefill and decode kernels are ❻ launched concurrently on the provisioned SMs. Bullet dynamically balancing the competing demands of TTFT and TPOT maintains high utilization with negligible overhead.
@@ 3.2. Performance Estimator
3.2.1. Problem Formulation
For a given LLM and hardware, the latency of co-executed prefill and decode is determined by six factors, termed as Execution State (): prefill sequence length (), batch size () and number of allocated SMs (), alongside decode context length (), batch size () and SMs (). Enumerating the millions of combinations for profiling is infeasible. Therefore, we build a lightweight analytical model with a minimal profile and runtime overhead. First, we propose SM-scaling roofline model (SRM) to derive single kernel latency under partitioned SMs without interference. Second, memory-subsystem contention is quantified when concurrent kernels are isolated in distinct SMs. Finally, the model is augmented with minimal sampled data for calibration. At runtime, the model executes within microseconds and continuously refines with online statistics.
3.2.2. SM-scaling Roofline Model (SRM)
We examine the performance of compute, memory access, and network communication when only of the SMs are available. Theoretical compute performance scales linearly as . Memory and network bandwidth exhibit proportional scaling until reaching inflection points and , where the SMs generate sufficient traffic to saturate the respective peak bandwidth and . Figure 8(a) illustrates the throughput of a memory-copy kernel, showing the inflection points of Nvidia A100 and H20 GPUs.
Given a kernel with operations and bytes of memory transactions, we construct an SM-scaling roofline model (SRM) in Equation 1 to estimate the theoretical latency on SMs. In the example of Figure 8(b), the memory bandwidth inflection point . When using 54 SMs, the attainable memory bandwidth remains at its peak, maintaining the original roofline slope while lowering the plateau. For 20 SMs, both the slope and plateau decline.
For each Execution State , we compute the arithmetic intensity of every LLM kernel with llm-viewer (Yuan et al., 2024), apply SRM to obtain its baseline latency on Np SMs, and aggregate these values to yield the total LLM latency . Since practical execution rarely matches the roofline bound, Equation 2 derives a scaling factor for calibrating SRM:
The factor is then extrapolated to unmeasured configurations, since the utilization pattern is near-linear between similar kernel inputs (Choi et al., 2022; Geoffrey et al., 2021; Chen et al., 2016). As shown in Figure 10, only two samples are sufficient to model the decode latencies by varying SMs. This demonstrates effective estimation aligned with kernel characteristics without extensive profiling.
![[Uncaptioned image]](x11.png)
3.2.3. Modeling Contention
When kernels execute on isolated SMs to prevent compute contention, memory subsystem and network contention persist. We verify that kernel latency remains stable under such interference, even if underlying hardware scheduling and inter-kernel resource competition vary. Extensive evaluation of 7360 concurrent on A100 (serve Llama3.1-8B (Dubey et al., 2024)) and 8H20 (serve Qwen3-32B (Yang et al., 2025)), each repeated for 30 runs, observes that 95% of latency measurements deviate less than 6.8% from the mean of the respective repeated runs. This tight distribution confirms the stability of kernel latency under varying prefill-decode interference patterns, which can be used as a reliable metric for contention modeling.
The worst-case memory subsystem interference can be quantified by co-executing a memory copy kernel on SMs and an up-gated layer (UG), which is a large-shaped GEMM in LLM, on the residual SMs. This is because prefill and decode kernels typically exhibit lower resource utilization than both UG and memory copy. Figure 10 shows that UG exhibits marginal performance degradation (¡8%) when using more than 60% SMs. Therefore, the latency of prefill kernels can be reliably calibrated via Equation 2 with minimal sample size. Conversely, memory-copy throughput decreases as UG sequence lengths grow. To model decode latencies, we measure the attainable bandwidth when concurrently executed with -length prefill to update SRM. Equation 2 is then applied for further refinement, since the memory interference pattern scales linearly with kernel utilization characteristics (Choi et al., 2022; Xu et al., 2018).
Inter-phase network contention is low since the traffic scales with either or , where is relatively small. The memory subsystem impact of network transfers is inherently lower than memory copy, and prefill-to-decode interference is already accounted for in the end-to-end calibration.
3.2.4. Profiling and Online Calibration
During offline profiling, Bullet records compute, memory, and network performance across varying SM counts in a single sweep to construct SRM, then measures prefill and decode latencies for a sparse set of values as calibration. Continuous online data collection for recalibration is straightforward. Leveraging only Equation 2 and linear extrapolation, the runtime overhead for update and prediction is negligible.
Table 2 shows that calibrated SRM maintains high accuracy even for inputs beyond the sampled range while keeping profiling overhead low. This precision is sufficient for reliable real-time scheduling optimization.
Phase (Data Range) | #Samples (Range) | MAPE |
18 (16k) | 9.1% | |
Prefill (32k) | 30 (32k) | 5.3% |
308 ( 16k, 128) | 10.5% | |
Decode ( 32k, 256) | 364 ( 32k, 256) | 7.7% |
@@ 3.3.1. Scheduling Workflow
Each of the prefill and decode concurrent execution engine (section 3.5) runs a scheduler autonomously. At every step, the scheduler reads system status from the global metadata buffer (section 3.5.1), forecasts latencies with performance estimator, and reorders pending requests. Upon detecting potential SLO violations, the scheduler greedily searches for the optimal configuration and invokes computational resource manager (section 3.4) to repartition SMs.
As illustrated in Figure 11, for prefill scheduler, a fixed number of layers is launched per step, and synchronizes for CPU (teal triangle) to make subsequent scheduling decisions. This enables fine-grained control over prefill progress temporally for rapid adaptation to system fluctuations. Conversely, decode scheduler issue kernels as a single CUDA Graph (Corporation, 2025) to eliminate the launch overhead of small kernels.
The primary scheduling objective is to prioritize prefill while respecting decode SLOs, since shorter prefill latency enlarges decode batch size and raises system throughput (Agrawal et al., 2023). During concurrent operation, the decode phase is provisioned with the minimum SM counts that satisfy SLO. As the final prefill layer approaches completion, additional SMs are allocated to decode phase to facilitate a smooth transition between co-running prefill/decode and decode-only. During reconfiguration, Bullet eliminates inter-phase synchronization by partially sharing SMs between phases rather than idling unused resources. Although a perfect non-overlapping allocation is challenging, layer-wise prefill scheduling confines such compute resource interference to minimal, predictable regions.
3.3.2. Request Scheduling and Resource Provisioning
To facilitate SLO-aware scheduling, Bullet monitors the system state defined as , where denotes the execution state (section 3.2.1), the prefill progress, and the per-request latency metrics. Prefill state comprises the queuing-request set , the in-flight request set , and the executed layer count . For every request , arrival time and decode-start time are recorded. At time , the prefill scheduler estimates GPU execution times for all requests in under current , while the decode scheduler predicts next-step latency and updates the corresponding TPOT. These estimations are then written to the global metadata buffer, allowing each scheduler perceive the holistic state.
Algorithm 1 outlines the prefill scheduler, and the decode scheduler follows analogously. The algorithm continuously monitors execution progress and updates latency estimates (lines 2-4). Requests in the waiting queue are reordered by ascending predicted latency (line 5) when the reordering does not violate TTFT SLOs for pending requests. This reduces average TTFT without starvation. To start a new prefill step (line 10-13), requests are batched until reaching arithmetic intensity limits under the current execution state. When provisioning resources, prefill phase is prioritized and provisioned with more SMs unless TPOT is compromised (line 15). In extreme cases of high request load, the decode phase would be temporarily suspended (Figure 11-➁) if TPOT SLO is still met. When both SLOs cannot simultaneously be satisfied, indicating the system is beyond maximum capacity, a balanced SM ratio is enforced to limit excessive latency in either phase (line 14).
@@ 3.4. Computational Resource Manager
While MPS (NVIDIA, 2025) with CUDA Green Context (gre, 2025) supports SM partitioning, its memory overhead exceeds 700MB for only 4 static policies in LLM serving (Cui et al., 2025), rendering it impractical for Bullet’s fine-grained control. For flexible low-overhead SM provisioning, we employ a library of SM masking techniques (Bakita and Anderson, 2023, 2024). The libsmctrl_set_stream_mask API sets the metadata of a CUDA stream (Corporation, 2025) to restrict all kernels launched within it to execute on a specified SM subset. Section 4.3.3 validates that this approach adds only microsecond-level runtime overhead and zero extra memory. Similar interfaces exist on other platforms, exemplified by AMD’s hipExtStreamCreateWithCUMask (hip, 2025).
Bullet creates a CUDA stream in each concurrent execution engine for prefill and decode. Whenever the scheduler issues a repartitioning command, the system immediately invokes the libsmctrl API to reconfigure the corresponding stream, thereby restricting all subsequent kernels to the newly provisioned SMs. This instantaneous configuration (Figure 11-➀➂), supports rapid adaptation to dynamic system states with flexible SM allocation.
@@ 3.5. Concurrent Execution Engine
3.5.1. Engine Architecture
Figure 12 illustrates the concurrent execution engines for prefill and decode, each residing in a separate process and driven by the corresponding scheduler. Both engines share a CPU buffer and unified GPU memory pool. The CPU buffer is implemented as OS-managed shared memory, stores global system states and employs compact control bits to indicate data availability, enabling low-latency status exchange. For GPU memory management, Bullet employs a dedicated initialization process that allocates model weights and KV cache (Kwon et al., 2023) prior to engine launch. The resulting memory region is shared between engines via cudaIpcGet/OpenMemHandle API (Corporation, 2025) with no adverse effects as documented. Equivalent facilities, such as AMD’s hipIpcOpenMemHandle (hip, 2025), allow the same design to be deployed on other hardware. Since the address space and metadata are shared, Bullet is fully compatible with existing KV cache and prefix cache optimizations. An atomic lock serializes allocation and deallocation transactions that may be issued concurrently by both engines, ensuring correctness with minimal performance impact.